40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 13 21:30:55 2020, shruti, Update, ALS, ALS OOL noise with arms locked 20200213_ALS.pdf
    Reply  Fri Feb 14 14:02:13 2020, shruti, Update, ALS, ALS OOL noise with arms locked 819472_10.pdf225642_10.pdf693500_10.pdf
       Reply  Fri Feb 14 14:52:41 2020, gautam, Update, ALS, ALS OOL noise with arms locked 
          Reply  Tue Feb 18 18:14:59 2020, shruti, Update, ALS, ALS OOL noise with arms locked ALSNoise.pdf
             Reply  Wed Feb 19 22:20:22 2020, rana, Update, ALS, ALS OOL noise with arms locked 
             Reply  Fri Feb 21 10:59:08 2020, shruti, Update, ALS, PDH error signals? IMG_0082.jpgIMG_0083.jpgIMG_0084.jpgIMG_0085.jpg
                Reply  Fri Feb 21 13:02:53 2020, Koji, Update, ALS, PDH error signals? 
                   Reply  Fri Feb 21 20:44:18 2020, shruti, Update, ALS, ALS OOL noise and PDH ALSNoise.pdfIMG_0086.jpg
                      Reply  Sun Feb 23 18:15:22 2020, rana, Update, ALS, ALS OOL noise and PDH 
Message ID: 15213     Entry time: Fri Feb 14 14:02:13 2020     In reply to: 15211     Reply to this: 15214
Author: shruti 
Type: Update 
Category: ALS 
Subject: ALS OOL noise with arms locked 

[Meenakshi, Shruti]

Even though we were not able to lock the the IR beat (by enabling LSC) during the day possibly because of increased seismic activity, we tried to the measure the ALS beat frequency noise by changing the PDH side-band frequency to different values.

I tried choosing values that corresponded to the peaks in the PM/AM as found in elog:15206 but for some reason unknown to us the cavity did not lock between 700-800 kHz.

The three attachments have data for different sideband frequencies:

Attachment 1: 819.472 kHz (peak in PM/AM, measured around noon)

Attachment 2: 225.642 kHz (peak in PM/AM, measured earlier in the morning)

Attachment 3: 693.500 kHz (not a peak in PM/AM)

We don't think these plots mean much and will do the measurement at some quieter time more systematically.

 

While doing the experiment, the ITMY pitch actuation was changed from -2.302 to -2.3172V because it locked better.

The ITMX, ETMX alignment was also tweaked to try to lock with different sideband frequencies and then restored to the values that were found earlier in the morning.

Attachment 1: 819472_10.pdf  72 kB  | Hide | Hide all
819472_10.pdf
Attachment 2: 225642_10.pdf  72 kB  | Hide | Hide all
225642_10.pdf
Attachment 3: 693500_10.pdf  72 kB  | Hide | Hide all
693500_10.pdf
ELOG V3.1.3-